# Efficient 3D Bayesian Full Waveform Inversion and Analysis of Prior Hypotheses

### v1

@@ 3.1 Inversion Results
Figures 1a - 1d display the inversion results obtained using the high frequency band data; results using low and intermediate frequency bands are provided in Appendix A. From left to right, each figure shows the true velocities, the posterior average velocity, standard deviation, and relative error maps, where the relative error is the absolute difference between true and mean velocities divided by the standard deviation at each point. The top two rows show vertical sections Y = 2.5 km and X = 2.5 km; bottom row shows a horizontal section Z = 1.25 km.
The three mean sections resemble the true velocity structures, and the overall relative errors are less than 3, indicating that differences between the true and inverted mean models are within three standard deviations, as would be expected of the true solution to this Bayesian problem.
@@
Figure 1: (a) - (d) High frequency 3D variational Bayesian FWI results obtained using PSVI and the Uniform prior distribution. (e) - (h) Inversion results after fine tuning the VPR results, as explained in the Discussion. In each figure, top two rows show vertical sections Y = 2.5 km and X = 2.5 km; bottom row shows a horizontal section Z = 1.25 km.
@@ 3.2 Testing Different Prior Hypotheses
In addition to the Uniform prior distribution, we introduce a set of smoothed prior distributions with different strengths of spatial smoothness controlled by a hyperparameter $\sigma_{s}$ (larger $\sigma_{s}$ allows rougher models - see Zhao and Curtis [12]). We consider 7 smoothed prior pdf’s with different $\sigma_{s}$ values: 2000, 1000, 500, 250, 125, 62.5 and 31.25, respectively. In the later part of this paper, we also denote the Uniform prior distribution as $\sigma_{s}=\infty$ (no smoothing). These 8 prior pdf’s are used to mimic a situation in which we have different prior hypotheses and wish to analyse their consequences.
@@
The posterior pdf’s using each of the 7 smoothed priors are obtained by replacing the Uniform prior pdf in inversion results from the previous section (Figures 1a to 1d) by the respective smoothed prior pdf’s, using VPR. Thus, no additional inversion is performed to obtain these results, and in Appendix B, we verify their quality by comparing them to results obtained by performing independent Bayesian inversions. Figure 2 displays the results using all 8 prior pdf’s on vertical section Y = 2.5 km, each row presenting the posterior pdf obtained using the $\sigma_{s}$ value indicated on the left. The posterior marginal pdf’s of two vertical profiles at locations X = 1 km and X = 3 km are displayed in Appendix C.
@@
Figure 2: Inversion results obtained using (a) the Uniform prior pdf, and (b) – (h) the 7 smoothed priors with decreasing values of $\sigma_{s}$ (increasing strength of smoothness), in vertical section Y = 2.5 km.
@@
As the magnitude of smoothness increases from Figures 2a to 2h, both the posterior samples and the mean velocity maps exhibit fewer spatial variations. The standard deviations decrease (note that colorbar scales vary in the third column); this leads to increased relative errors wherever (true) large velocity contrasts between neighbouring cells are precluded by the smoothing prior information for small $\sigma_{s}$.
@@
By comparing each set of results with the true velocity model in Figure 1a, we find that smoothed prior distributions with $\sigma_{s}=1000$ or $\sigma_{s}=500$ (Figures 2c and 2d) are relatively good choices for this FWI problem. However, in reality we never know the true velocity structure, so can not calculate or use relative errors to choose between prior distributions. In deterministic FWI, an order of preference is often constructed by calculating data misfit values arising from the different inversion results (referred to as an L-curve). Following a similar approach, we compare the 8 posterior distributions by drawing N = 100 samples from each posterior pdf and performing forward simulations to calculate corresponding misfit values between the observed and synthetic data. This results in 8 approximate probability distributions over misfit values from the inversion results using different priors, displayed in Figure 3a. As expected, the overall misfit values increase as more smoothness is injected into the inversion.
The mean and standard deviation of each distribution are displayed in Figure 3b by black dots and error bars, and are connected by a red line which we call a Bayesian L-curve. From left to right along the curve, the magnitude of prior information (relative smoothness in this case) decreases from $\sigma_{s}=31.25$ (top-left point) to $\sigma_{s}=\infty$ (no smoothing – bottom-right point). An arrow marks the case with $\sigma_{s}=1000$, and we observe that data misfits start to decrease significantly if we further increase the roughness value $\sigma_{s}$ of the prior information, implying that this value provides a balance between information from the smoothing distribution and the observed data.
More generally, many different types of prior hypotheses can be represented by a hypothesis space $\mathcal{H}$ represented schematically in Figure 3c. Given a rather non-informative Uniform prior distribution in the old posterior pdf, in principle we can inject any additional prior assumptions into the inversion; each would have different effects on the inversion results, between which we may wish to discriminate. As an example, we consider another prior assumption using the real-geology informed prior distribution from Zhao and Curtis [12] - described in Appendix D. The dashed blue line in Figure 3b displays the corresponding Bayesian L-curve for this set of prior assumptions, where the strength of geological prior information decreases from left to right. The data misfits are lower than those from smoothing priors in red because geologically informed priors impose less smoothness vertically than horizontally (which is correct in the true model - and often in the Earth).
@@
Figure 3: (a) Probability density functions for data misfit values calculated from posterior pdf’s displayed in Figure 2 using differently smoothed priors. (b) Bayesian L-curves representing average data misfits and their standard deviations (error bars) with respect to different prior hypotheses. The horizontal axis represents the normalised (relative) weakness of prior information applied by each prior distribution: the value 1 represents the Uniform prior pdf. The red line denotes the L-curve using smoothed prior pdf’s, and dashed blue line denotes the equivalent curve when adding the geological prior information. (c) A space $\mathcal{H}$ containing all possible prior hypotheses. Black dot represents a non-informative Uniform prior distribution, and we denote different prior assumptions such as smoothed and geological prior pdf’s in this study by different paths through $\mathcal{H}$.
@@ 3.3 Computational Cost
The inversion using the Uniform prior distribution (Figures 1a to 1d) is obtained by performing 5200 forward and adjoint simulations. To further reduce the computational cost, we use a minibatch of 36 shot gathers, randomly selected from the total of 81 shots in each FWI simulation [11]. Note that in deterministic FWI, one would usually perform $\sim$100 iterations to obtain a reasonable result, and a full batch of 81 shots is often used at each iteration. Since forward simulation using 81 shot gathers would require 2.25 times more computation than that using a minibatch of 36 shots, the computational cost of 100 iterations implemented in deterministic FWI is equivalent to 225 forward evaluations performed in this study.
With PSVI we therefore obtain reasonable probabilistic uncertainty estimates using only an order of magnitude more computation than deterministic FWI, from which no robust uncertainty information can be obtained. In addition, this single set of variational inversion results allows multiple posterior distributions corresponding to different prior pdf’s to be produced using VPR, without any further FWI simulations [12].
@@ 4 Discussion
Rather than approximating the true solution to the desired inverse problem directly, the VPR approximation to each new posterior pdf includes only information about the true solution that is already included within in the old posterior pdf. It therefore carries over an imprint of any errors in the latter [12]. VPR results may therefore differ from independent Bayesian inversion results. To improve the results, we can refine (fine tune) the outcomes obtained from VPR using observed data. We do this by performing another variational Bayesian inversion using PSVI, but with a smaller number of iterations which start from the VPR output, because the VPR solution should already be reasonably close to the true posterior pdf. Figures 1e to 1g show the fine-tuned inversion results with 800 additional simulations, which are improved compared to the VPR results (Figure 7 in Appendix B). To reduce this additional computation we might again depart from formal Bayesian analysis by using the Bayesian L-curve obtained from VPR to select a smaller subset of prior hypotheses to entertain, then fine tune only those VPR posterior pdf’s.

### v2

@@ 3.1 Comparison with Previous Results
Figure 2: Single frequency band inversion results obtained using 4 different methods, at vertical section Y = 2.5 km. The true velocity model is displayed in Figure 1c. From top to bottom, each row shows the posterior mean velocity, standard deviation (std), and the relative error maps, respectively, where the relative error is the absolute difference between true and mean velocities (the Error) divided by the standard deviation at each point. Dashed black boxes highlight structures that are not inverted correctly by any method due to cycle skipping.
To justify the above statements and to better compare the 4 sets of results, in Figure 3 we compare the posterior marginal distributions obtained from the 4 methods along two vertical profiles at horizontal locations of 1 km (top row) and 3 km (bottom row), respectively. The locations of these two profiles are displayed by dashed red lines in Figure 1c. In Figure 3, red lines show the true velocity profile. Overall, the marginal distributions in Figures 3c and 3d present similar features and are broader than those in Figures 3a and 3b, indicating that both PSVI and sSVGD provide better posterior uncertainty estimates on average compared to the other two methods.
Figure 3: Posterior marginal pdf’s obtained using different methods, extracted at two locations marked by red dashed lines in Figure 1c. Red lines show the true velocity profile. White arrows highlight posterior marginal pdf’s that present incorrect phase information caused by cycle skipping.
However, the above 4 results, especially those obtained using PSVI and sSVGD, show some consistent but incorrect patterns. First, the posterior marginals in the first row in Figure 3 provide biased phase information in which the true velocity profile is excluded from the posterior pdf’s, as marked by white arrows. PSVI uncertainties are significantly narrower in that section of the profile accentuating the error, presumably because the algorithm is stuck in a local minimum of the variational optimisation problem that happens to have low values for the velocity standard deviations. This section corresponds to the results highlighted by the dashed black boxes in Figure 2, due to the incorrectly inverted low-wavenumber component (cycle skipping). Second, in Figure 3 we observe roughly the same magnitude of posterior uncertainties (i.e., similar width of those marginal pdf’s) in both shallower and deeper parts of the model. Given that the sensitivity of surface seismic data should often be higher in the near surface than in deeper parts of the Earth, we would expect to observe higher uncertainties at depth. This is caused by additional prior information (spatial smoothness) imposed into the inversion.
In the following section, we solve these two issues to improve the inversion results.
3.2 Improved Inversion Results
In this section, we show how to improve the 3D FWI results. To avoid cycle skipping, similarly to the common strategy in linearised FWI problems, we employ a multiscale inversion approach by inverting waveform data from low frequency to high frequency [10]. We consider three frequency bands referred to as the low (1 – 6 Hz), intermediate (1 – 10 Hz), and high frequency band (1 – 13 Hz). Variational optimizations using data in intermediate and high frequency bands used results from the previous band as a starting point. The starting point for the low frequency band variational inversion is set to be a standard Gaussian distribution within a unconstrained space, as shown in equation 3 and suggested by [22]. In preparation for subsequent sections, we refer to the final solution using high frequency data as $q_{old}(\mathbf{m})$.
Any regularisation (a form of prior information) is somewhat arbitrary as we will never know whether the imposed prior information is consistent with the true Earth or not a priori, and thus whether the regularisation will spuriously bias the inversion results or not. We therefore remove all regularisations (note that spatial smoothing is imposed in the previous tests), and simply use the uniform pdf displayed in Figure 1b as prior information for FWI. As long as the posterior solution is not unduly constrained by the uniform distribution bounds, we assume that the inversion is constrained mainly by the likelihood function and hence by observed waveform data. Note that this will increase the effective support of the posterior pdf (the hyper-volume over which it is significantly greater than zero), and thus will also increase the complexity of this FWI problem, compared to more strongly regularised problems (e.g., with tighter uniform bounds). Due to limited computational resources, we only perform PSVI in this test. After obtaining reasonable inversion results using this non-informative, broad prior distribution, in the next section we analyse different prior assumptions efficiently using a variational prior replacement methodology.
Figure 4 displays the final inversion results obtained using the high frequency band data; interim results using low and intermediate frequency bands are presented in Appendix A. The top two rows display two vertical sections at Y = 2.5 km and X = 2.5 km. The bottom row illustrates one horizontal section at a depth Z = 1.25 km. The three sections are marked by dashed black lines in Figure 1a. From left to right, Figures 4a – 4d show the three true velocity sections, the average velocity, the standard deviation, and the relative error maps of the posterior probability distribution, where the relative error is the absolute difference between the true and mean velocities divided by the standard deviation at each point.
@@ 3.2 Improved Inversion Results
Figure 4: High frequency 3D variational Bayesian FWI results obtained using PSVI and the uniform prior distribution displayed in Figure 1b. Top two rows show vertical sections at Y = 2.5 km and X = 2.5 km; the bottom row shows a horizontal section at Z = 1.25 km. Panels (a) – (d) display the true velocity model, the mean, the standard deviation and the relative error (calculated by the absolute difference between the true and mean velocities divided by the standard deviation at each point) maps of the posterior distribution, respectively.
The three mean sections resemble the true velocity structures (note that even for an entirely correct solution, this need not be the case in probabilistic inversions since the mean of the models is a statistic, not a model in itself). This is especially true in the shallow subsurface above 2 km where high resolution structures are accurately inverted. However, the mean and true values diverge at deeper levels, possibly due to reduced data sensitivity since the standard deviations broadly increase with depth. The overall relative errors are less than 3, indicating that differences between the true and inverted mean models are within three standard deviations, as would be expected of the true solution to this Bayesian problem. Note that in Figure 4d, higher relative errors (shown in yellow) are present near the surface because in fact the true model deliberately lies outside of the prior bounds shown in Figure 1b, to assess the method’s behaviour in such circumstances [50]. Overall, the inverted images and the posterior uncertainties in Figure 4 are significantly more accurate than those displayed in Figure 2.
@@ 3.3 Testing Different Prior Hypotheses
In addition to the uniform prior distribution, we now consider a set of smoothed prior distributions. Define a second-order finite difference operator $\mathbf{S}$ to calculate curvature of model parameter values between adjacent grid cells in a spatial 3D velocity model represented by model vector $\mathbf{m}$ [53]. This is further applied to $\mathbf{m}$ to impose spatial smoothness, by imposing a Gaussian distribution on the product $\mathbf{Sm}$
$p(\mathbf{Sm})\propto\exp\left(-\frac{(\mathbf{Sm})^{T}(\mathbf{Sm})}{2\sigma_{s}^{2}}\right),$ (6)
where $\sigma_{s}$ is a hyperparameter that controls the strength of spatial smoothness assumed (a smaller $\sigma_{s}$ causes the probability of any particular rough model to become lower). Equation 6 can be interpreted as applying a Tikhonov matrix $\mathbf{S}$ to $\mathbf{m}$ [18]. Then the smoothed model prior pdf can be expressed as
$p_{smooth}(\mathbf{m})\propto p(\mathbf{Sm})p_{old}(\mathbf{m}),$ (7)
@@
where $p_{old}(\mathbf{m})$ is the (old) uniform prior distribution defined in the previous section. We consider 7 smoothed prior pdf’s with different $\sigma_{s}$ values: 2000, 1000, 500, 250, 125, 62.5 and 31.25, respectively. In the later part of this paper, we also denote the uniform prior distribution as $\sigma_{s}=\infty$ indicating no smoothing. These 8 prior pdf’s are used to mimic a real situation in which we have different prior hypotheses and wish to analyse their consequences.
@@
The posterior pdf’s using each of the 7 smoothed priors are obtained by replacing the uniform prior pdf in inversion results $p_{old}(\mathbf{m}|\mathbf{d}_{obs})$ from the previous section (Figure 4) by the respective smoothed prior pdf’s, using variational prior replacement (VPR). Thus, no additional inversion is performed to obtain these results. In Appendix B, we verify the results of VPR by comparing them to those obtained by performing independent Bayesian inversions (verification of VPR in 2D FWI was also conducted in Zhao and Curtis [53]). Figure 5 displays the results using all 8 prior pdf’s on vertical section Y = 2.5 km. From left to right, each column shows one posterior sample, the mean, the standard deviation, and the relative error maps of the posterior distribution, respectively. From top to bottom, each row presents the posterior pdf obtained using the $\sigma_{s}$ value indicated on the left of the figure.
@@
Figure 5: Inversion results obtained using (a) the uniform prior pdf and (b) – (h) the 7 smoothed priors with decreasing values of $\sigma_{s}$ (equivalently increasing smoothness), in a vertical section Y = 2.5 km. From left to right, they represent one posterior sample, the mean velocity, the standard deviation and the relative error maps.
@@
From Figures 5a to 5h, as the magnitude of smoothness increases, stronger prior information favouring spatially smooth structures is injected into the inversion results. Consequently, both the posterior samples and the mean velocity maps become smoother with fewer spatial variations. The standard deviations decrease (note that different colorbar scales are used in the standard deviation maps in the third column); this leads to increased relative errors, as (true) large velocity contrasts between neighbouring cells are precluded by the prior information for small $\sigma_{s}$.
Figure 6 compares the posterior marginal pdf’s of two vertical velocity profiles at locations X = 1km (top row of Figure 6) and X = 3km (bottom row), marked by red dashed lines in Figure 1c. Figure 6a displays the posterior marginal pdf’s obtained from the uniform prior distribution, and Figures 6b – 6h show those from the smoothed prior pdf’s – the smoothness values $\sigma_{s}$ are noted in the title. In each figure, the red line shows the true velocity profile and black line displays the inverted mean velocity profile. We observe that with the increase of the magnitude of smoothness injected by the prior information, the posterior marginal pdf’s become smoother with fewer spatial variations (thus lower uncertainties), similarly to the conclusions drawn above.
Figure 6: Posterior marginal pdf’s at two vertical locations marked by red dashed lines in Figure 1c. (a) Results obtained using the uniform prior pdf, and (b) – (h) results using the smoothed prior pdf’s with different $\sigma_{s}$ values.
By comparing each set of results with the true velocity model in Figure 1c, we find that the posterior pdf’s are more tightly constrained around the true values as we increase the magnitude of the smoothness up to $\sigma_{s}=500$ in Figure 5d, especially in deeper parts of the model. Interestingly, in the third column in Figures 5c and 5d, we observe higher uncertainties at the boundaries between strata below 1.5 km depth, capturing the expected ’uncertainty-loop’ character of posterior uncertainties proposed by Galetti et al. [15]. However, Figure 5 shows that uncertainty loops appear because smoothness is applied to models, and do not appear when models are allowed to be rough; this is consistent with the results of Galetti et al. [15] which were obtained using a trans-dimensional Monte Carlo travel-time tomography method that imposed strict smoothness (indeed, constant velocity) over large areas of model. We therefore qualify the proposition of Galetti et al. [15], by hypothesising that posterior uncertainty loops are in fact a product of smoothness constraints in tomographic problems (including FWI).
The inversion results become worse in Figures 5e – 5h, as the effect of prior information overpowers information in the observed data, resulting in biased posterior solutions and overly smoothed structures. Not surprisingly, most of the relative errors are then larger than 3. We conclude that the smoothed prior distributions with $\sigma_{s}=1000$ or $\sigma_{s}=500$ (Figures 5c and 5d) are relatively good choices for this FWI problem.
@@
However, in reality we never know the true velocity structure, so can not calculate or use relative errors to choose between prior distributions, other than in synthetic tests such as in Figures 5 and 6. In deterministic FWI, an order of preference is often constructed by calculating data misfit values arising from the different inversion results (referred to as an L-curve). Following a similar approach, we compare the 8 posterior distributions by drawing N = 100 samples from each posterior pdf and performing forward simulations to calculate corresponding misfit values between the observed and synthetic data. This results in 8 approximate probability distributions over misfit values from the inversion results using different priors, displayed in Figure 7a. As expected, the overall misfit values increase as more smoothness is injected into the inversion.
@@
Figure 7: (a) Probability density functions for data misfit values calculated from posterior pdf’s displayed in Figure 5 using differently smoothed priors. (b) Bayesian L-curves representing the average data misfits and their standard deviations (error bars) with respect to different prior hypotheses. The horizontal axis represents the normalised (relative) weakness of prior information applied by each prior distribution: the value 1 represents the non-informative, uniform prior pdf represented in Figure 1b. The red line denotes the L-curve found by requiring smoothness in the prior pdf’s, and dashed blue line denotes the equivalent curve when requiring consistency with geological prior information. (c) A schematic diagram of the space $\mathcal{H}$ containing all possible prior hypotheses. Black dot represents a non-informative uniform prior distribution, and we denote different prior assumptions such as smoothed and geological prior pdf’s in this study by different paths (arrows) through $\mathcal{H}$.
The mean and standard deviation of each distribution are displayed in Figure 7b by black dots and red error bars, and are connected by a red line which we call a Bayesian L-curve. From left to right along the curve, the magnitude of prior information (relative smoothness in this case) decreases from $\sigma_{s}=31.25$ (top-left point) to $\sigma_{s}=\infty$ (no smoothing – bottom-right point). An arrow marks the case with $\sigma_{s}=1000$, and we observe that data misfits start to decrease significantly if we further increase the roughness value $\sigma_{s}$ of the prior information beyond this, implying that this value provides a balance between information from the smoothing distribution (used to narrow the possible parameter space) and the observed data (used to achieve good data fits).
From a Bayesian point of view, if we do not know the smoothness of the true Earth, then we have little reason to choose any one of these posterior distributions – they are all part of the posterior uncertainty. We may have sufficient prior intuition to suspect that posterior models are too rough in Figure 5a and too smooth in Figure 5h, but it is difficult to establish a probability distribution over levels of smoothness a priori. Hierarchical Bayesian methods might be used to estimate this distribution a posteriori. However, such an approach introduces physical inconsistency to Bayesian inversion results [27]. Alternatively, in a departure from Bayesian methods we might decide to entertain only a subset of these various results, for example those around $\sigma_{s}=1000$.
More generally, many different types of prior hypotheses can be represented by a hypothesis space $\mathcal{H}$ represented schematically in Figure 7c. Given a rather non-informative uniform prior distribution in the old posterior pdf, in principle we can inject any additional prior assumptions into the inversion; each would have different effects on the inversion results, between which we may wish to discriminate. As an example, we consider another prior assumption from real-geology. The geology informed prior information is obtained by subsampling a set of realistic geological images using a predefined window size. These subimages are used to calculate a local correlation matrix between parameter pairs that lie within that window size to define prior correlations between pairs of model parameters used in this test [53]. We consider 5 such geological prior distributions constructed using different local cubic windows of edge length 2, 4, 6, 8 and 10 cells, respectively. Again, the corresponding posterior distributions are calculated using VPR by replacing the old uniform prior distribution by each geological prior distribution in turn.
Figure 8 displays the corresponding posterior pdf’s, where the sizes of the correlation windows used are denoted on the left side of Figure 8. Figure 8a represents the results obtained using the uniform prior distribution. We observe that larger prior correlation windows generally result in better inversion results that are more similar to the true velocity model (Figure 1c), since a larger window can define better prior correlation information, which improves the inversion results. However, note that a larger window also has higher memory requirements.
These inversion results are then used to construct a Bayesian L-curve, as displayed by the dashed blue line in Figure 7b. From left to right along this dashed blue line, the correlation window size decreases from $10\times 10\times 10$ (top-left point) to $0\times 0\times 0$ (bottom-right point), the latter being the uniform prior pdf with no prior correlation. This indicates that the strength of geological prior information decreases from left to right. The data misfits are lower than those from smoothing priors in red because geologically informed priors impose less smoothness vertically than horizontally (which is correct in the true model - and often in the Earth).
@@ 3.4 Computational Cost
The inversion using the uniform prior distribution (Figure 4) is performed within three frequency bands, each involving 200, 300 and 800 iterations for variational inference (optimisation) in the low, intermediate and high frequency bands respectively. In each iteration, 4 random samples are used to estimate expectations in the EBLO[$q(\mathbf{m})$] in equation 2. Therefore, a total of 5200 samples, thus 5200 forward and adjoint simulations, are used to obtain the inversion results displayed in Figure 4. To further reduce the computational cost, we use a minibatch of 36 shot gathers, randomly selected from the total of 81 shots in each FWI simulation [50]. For comparison, the same test conducted in Zhang et al. [50] were obtained using 4000, 400,000 and 80,000 samples (forward and adjoint simulations) when using automatic differentiation variational inference ADVI [ADVI – 22], Stein variational gradient descent [SVGD – 25], and stochastic SVGD [sSVGD – 16], respectively. SVGD and sSVGD are far more expensive than PSVI; ADVI is computationally cheaper but provides biased inversion results with strongly underestimated posterior uncertainties [50].
Note that in linearised, deterministic FWI, one would usually perform $\sim$100 iterations to obtain a reasonable result, and a full batch of 81 shots is normally used at each iteration to provide accurate gradient information. Since forward simulation using 81 shot gathers would require 2.25 times more computation than that using a minibatch of 36 shots, the computational cost of 100 iterations implemented in deterministic FWI is equivalent to 225 forward evaluations performed in this study.
With PSVI we therefore obtain reasonable probabilistic uncertainty estimates using only an order of magnitude more computation than deterministic FWI from which no robust uncertainty information can be obtained. In addition, this single set of variational inversion results allows multiple posterior distributions corresponding to different prior pdf’s to be produced using VPR, without any further FWI simulations even in 3D.
@@ 4 Discussion
All of the posterior pdf’s with different priors were obtained from only a single Bayesian 3D FWI using a uniform prior distribution. However, to translate uncertainties from model parameter space into data space, hundreds of additional forward simulations were performed to obtain data misfit values to calculate the Bayesian L-curves displayed in Figure 7. Alternatively, boosting variational inference (BVI) implicitly provides a small number (tens) of representative samples that represent major components of the uncertainties [51]. Those representative samples might be used to construct Bayesian L-curves at a significantly reduced computational cost compared to the approach used here.
A common way to discriminate between prior hypotheses in Bayesian inference is to calculate the evidence term,
[…]
for each hypothesis, which requires an integral over the entire model parameter space to be evaluated. While this is computationally possible for some low dimensional problems [36], for Bayesian 3D FWI it is infeasible due to the high dimensionality of the problem. Hence, the approach used here is the first to allow such models and prior hypotheses to be compared in a probabilistic manner and at feasible computational cost in 3D.
While we used prior distributions with varying degrees of smoothness to simulate different prior hypotheses, more realistic prior distributions that encode geological information can be employed in Bayesian inversion [28, 24, 9, 38]. By using the VPR framework we can calculate the corresponding posterior distributions and analyse different prior hypotheses at relatively low cost.
Rather than approximating the true solution to the desired inverse problem directly, the VPR approximation to each new posterior pdf includes only information about the true solution that is already included within in the old posterior pdf. It therefore carries over an imprint of any errors in the latter [53]. VPR results may therefore differ from independent Bayesian inversion results. This effect is potentially non-negligible due to the higher dimensionality and complexity of 3D Bayesian FWI problems as shown in Appendix B; qualitatively the results are less good than those obtained for 2D FWI in Zhao and Curtis [53]. To improve the results, we can use a relatively lower cost calculation to refine (fine tune) the outcomes obtained from VPR using observed data. We do this by invoking Bayes’ rule again (performing another variational Bayesian inversion using PSVI), but with a smaller number of iterations which start from the VPR output because the VPR solution should already be reasonably close to the true posterior pdf. For example, we fine tuned the above VPR results by performing variational inversion with an additional 200 iterations using 4 samples per iteration. This costs an extra 800 forward and adjoint simulations, and the results are displayed in Figure 9. These are improved compared to the initial VPR results (Figure 12) and are more similar to independent inversion results (Figure 13). If we wanted to reduce this additional computation further, we might again depart from formal Bayesian analysis by using the Bayesian L-curve obtained from VPR to select a smaller subset of prior hypotheses to entertain, then fine tune only those VPR posterior pdf’s.
Figure 9: Inversion results after refining the VPR results, which are more similar to independent variational Bayesian inversion results (presented in Appendix B).